Grade upto college level Physical Chemistry

The standard enthalpy of combustion at 25oC of hydrogen, cyclohexene (C6H10) and cyclohexane (C6H12) are -241 -3800 and -3920 kJ/mole respectively. Calculate the heat of hydrogenation of cyclohexene.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
The required reaction is
C6H10(g) + H2(g) → C2H12,(g), ∆H1 = ? ….(1)
Cyclohexene Cyclohexane
The given facts can be written as :
H2(g) + ½O2(g) → H2O, ∆H2 = - 241 kJ/mol …(2)
C2H10(g) + 17/2 O­2 → 6CO2, (g) + 5H2O; ∆H3 = - 3800 kJ/mol …(3)
C6H10(g) + 902(g) → 6CO2(g) + 6H2O, ∆H4 = - 3920 kJ/mol ….(4)
The required reaction (1) can be obtained by adding equations (2) and (3) and subtracing (4) from the sum of (2) and (3).
C2H10(g) + H2(g) → C6H12(g)
∆H1 = (∆H2 + ∆H3) -∆H4
= [-241 + (-3800)] – (-3920)
= (-241 - 3800) – (- 3920)
= - 4041 + 3920 = - 121 kJ/mole

Thanks
jitender
askIITians Faculty